Find the marginal rate of technical substitution of Q=ln L + ln K
1
Expert's answer
2014-10-21T12:33:16-0400
Q = ln L + ln K
In microeconomic theory, the Marginal Rate of Technical Substitution (MRTS) is the amount by which the quantity of one input has to be reduced when one extra unit of another input is used, so that output remains constant.
MRTS = MPL/MPK = Q(L)'/Q(K)' = (1/L)/(1/K) = K/L